OpenStudy (anonymous):

Here's a somewhat "problem-solving" type question I was doing earlier on today in a written math contest. I don't recall the question in verbatim but remember it pretty well. Whoever solves first gets a medal :D "Find all linear functions \(\bf f(x)=ax+b\), if \(\bf g(x) = f^{-1}(x)\), then \(\bf f(x) - g(x) = 44\)."

OpenStudy (zzr0ck3r):

well f(x) = ax+b is only linear if b = 0:) that should say find all affine functions

OpenStudy (anonymous):

@zzr0ck3r Remember I don't recall the question in verbatim. Also, the wording of the question implies that f(x) is a linear function hence a and b are constants...Also the question implies the invertibility of f(x) hence a cannot be 0. I shouldn't be telling you this but whatever lol :D

OpenStudy (zzr0ck3r):

was more a joke. Im sure it says linear...very common thing.

OpenStudy (zzr0ck3r):

b cant be constant

OpenStudy (zzr0ck3r):

unless its 0 linear must go through the origin

OpenStudy (anonymous):

@zzr0ck3r when I say "constant", I don't mean that b is fixed...

OpenStudy (anonymous):

this is a trick. he is taking a test and doesnt want it to bee obvious. also, he is an retriceole.

OpenStudy (zzr0ck3r):

No he aint

OpenStudy (anonymous):

retricehole

OpenStudy (anonymous):

A S S H O L E

OpenStudy (anonymous):

@tylerj37 I already solved this problem this morning before going to bed. just wanted to see who solves it first here on openstudy. I did it in about 5-6 minutes I think...with the writing and stuff combined

OpenStudy (zzr0ck3r):

Im saying @genius12 that y=mx+b is only linear if b=0. A linear equation must go through the origin.

OpenStudy (agent0smith):

y = ax + b so the inverse is y = (x-b)/a \[\Large ax + b -\frac{ x-b }{ a }= 44\]

OpenStudy (zzr0ck3r):

most people call affine equations linear when they are not.

OpenStudy (anonymous):

@zzr0ck3r The question says "Linear function". In the current context, a linear function is defined to be a polynomial of degree 0 or 1. Hence b is not required to be 0.

OpenStudy (anonymous):

@agent0smith That would be the first step. Keep going.

OpenStudy (zzr0ck3r):

well that's wrong:) a linear function is what it is...there are rules:)

OpenStudy (anonymous):

@zzr0ck3r You're wrong. Sorry to say.

OpenStudy (zzr0ck3r):

go read man....

OpenStudy (zzr0ck3r):

do you know what a linear transformation is?

OpenStudy (zzr0ck3r):

how then can mx+b be a linear transformation?

OpenStudy (zzr0ck3r):

its not important...again I was just making a joke as it is a very common misconception

OpenStudy (anonymous):

@zzr0ck3r You are referring to linear algebra. A linear function - a linear map - is different from a linear function in calculus and related areas.

OpenStudy (zzr0ck3r):

no its not

OpenStudy (anonymous):

yes it is go read

OpenStudy (zzr0ck3r):

I don't need to read, my analysis teacher had this discussion in front of my class 3 days ago

OpenStudy (zzr0ck3r):

show me anything that says what you are saying....

OpenStudy (anonymous):

That's great lol. Doesn't prove you right.

OpenStudy (zzr0ck3r):

again, one thing.

OpenStudy (zzr0ck3r):

just one

OpenStudy (zzr0ck3r):

:)

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Linear_function

OpenStudy (anonymous):

now please don't say that wikipedia isn't a scholarly source lol...

OpenStudy (zzr0ck3r):

lol

OpenStudy (zzr0ck3r):

ok man

OpenStudy (zzr0ck3r):

its the definition of linear

OpenStudy (zzr0ck3r):

why would it be different for calculus...

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

there is lots of terms used differently in various fields of mathematics. not each term is defined the same way in every field.

OpenStudy (zzr0ck3r):

again very common misconception...

OpenStudy (anonymous):

ok you need to stop with the misconceptions. lol. this is ain't no misconception. =] Also that link doesn't help prove your argument..

OpenStudy (zzr0ck3r):

you cant prove notation, it is what it is by definition

OpenStudy (zzr0ck3r):

you give me the definition that wiki gives you, ill stick to my analysis book.

OpenStudy (anonymous):

like i said, what subject/field is your "analysis" book for

OpenStudy (zzr0ck3r):

pick one

OpenStudy (zzr0ck3r):

dude I just showed you a site that says common mistakes....and then goes on to list the exact thing im talking about first. I know its not proof, but you might want to go read more than wiki about it.

OpenStudy (zzr0ck3r):

again, was not trying to start anything. just something I say when people say linear.

OpenStudy (zzr0ck3r):

sort of like when people say derive when they mean differentiate. I cant help it.

OpenStudy (agent0smith):

@zzr0ck3r if you were making a joke, how about not sh*tting up this question with it? Start a new question on the definition of a linear function and go nuts there. \[\Large x(a-\frac{ 1 }{ a}) + b + \frac{ b }{ a } = 44\] a-1/a = 0, and b+b/a =44

OpenStudy (zzr0ck3r):

ps. calculus is mostly performed on vector spaces, there is no seperation

OpenStudy (zzr0ck3r):

like the real numbers...

OpenStudy (agent0smith):

a =1 b = 22

OpenStudy (agent0smith):

I can post more working, but the thread's already a mess.

OpenStudy (zzr0ck3r):

we are smart, we can read around comments.

OpenStudy (anonymous):

@zzr0ck3r The author of that page calls linear functions in calculus as affine functions and says that the term "linear functions" should be defined the way it is liner algebra to avoid confusion. Once again, he's just saying that to avoid confusion. That in no way however proves to say that a linear function cannot be defined the way it is in calculus. Now in the future, a math conference may conclude that the term "linear function" will be replaced with the term "affine function" but for now, we are sticking to how a linear function is defined in the current context. @agent0smith Your answer is correct. But how many such f(x) that satisfy the conditions are there then?

OpenStudy (agent0smith):

Only one that i found, f(x) = x + 22 a=-1 isn't a valid solution.

OpenStudy (zzr0ck3r):

again you are acting like linear algebra and calculus are independent of each other.

OpenStudy (anonymous):

@agent0smith Good job. There is a single solution y = x + 22. That's end of this thread. Cheers.

OpenStudy (zzr0ck3r):

lol